OpenStudy (anonymous):

What is the slope intercept form of the equation x + 3y = -6?

OpenStudy (uri):

y=mx+b

OpenStudy (anonymous):

y=mx+c use y be the subject first

OpenStudy (uri):

3y=-x-6

OpenStudy (uri):

solve further

OpenStudy (anonymous):

I think it's y=3x+6

OpenStudy (uri):

ah y=-1/3x-6/3 y=-1/3-2

OpenStudy (anonymous):

Ohhh okay